solve the equation in the complex number system:

45x^2+1=6x

Answer by dabanfield(803)

45x^2+1=6x
45x^2 - 6x + 1 = 0
Using the quadratic formula with a = 45, b = -6 and c = 1 we have the two solutions:
x = (-(-6) + sqrt((-6)^2 - 4*45*1))/(2*45) =
(6 +(sqrt(36-180))/90 =
(6 + sqrt(-144))/90 =
(6 + sqrt(-1)*(sqrt(144))/90
(6 + i*12)/90 =
(6 + 12i)/90
and
x = (-(-6) - sqrt((-6)^2 - 4*45*1))/(2*45) =
(6 - (sqrt(36-180))/90 =
(6 - sqrt(-144))/90 =
(6 - 12i)/90